# Pauli blocking: probing beyond-mean-field effects in neutrino flavor evolution

### v1

@@ VI Outlook
The physics of flavor conversion due to neutrino-neutrino interactions shows a very rich phenomenology that we are still far from fully grasping. Most of the existing work on this topic is carried out employing the mean-field approximation. In this paper, we have attempted to investigate whether beyond-mean-field effects on neutrino-neutrino coherent forward scattering could introduce flavor instabilities that are not otherwise foreseen. To this purpose, we have focused on the regime where neutrinos are degenerate and introduced a heuristic prescription for Pauli-blocking terms in the neutrino self-interaction Hamiltonian. Such corrections define an effective ELN distribution, which modifies the (anti)neutrino dispersion relation. Hence, ensembles that are on the verge of stability in the mean-field case can become unstable when the beyond-mean-field terms are included and vice versa.
Relying on the gyroscopic pendulum analogy introduced for fast flavor conversion [16], we have investigated how the non-linear regime of the flavor evolution is modified by our beyond-mean-field prescription. Within the mean-field approximation, each (anti)neutrino polarization vector precesses around the potential generated by all other neutrinos in the ensemble. In the beyond-mean-field scenario, additionally, the length of the polarization vectors changes due to dynamical decoherence. The terms that cause this effect introduce an energy dependence in the collective potential and break the symmetry between the neutrino and the antineutrino equations of motion. As a result, flavor instabilities cascade down to small angular scales, and the evolution of the system is not periodic anymore.
Qualitatively, the flavor conversion outcome including the beyond-mean-field effects in our toy model is similar to what observed when incoherent scattering and vacuum mixing are included in the mean-field equations of motion, although the physics linked to each of these phenomena and their interplay with the self-interaction term in the Hamiltonian is different. We have neglected these additional terms in the equations of motion throughout this work, to highlight the impact of beyond-mean-field corrections on fast instabilities. However, collisions, vacuum mixing, and potential beyond-mean-field corrections (due to momentum correlations caused by neutrino degeneracy) should coexist in the core of an astrophysical system.

### v2

@@ VI Outlook
The physics of flavor conversion due to neutrino-neutrino interactions shows a very rich phenomenology that we are still far from fully grasping. Most of the existing work on this topic is carried out employing the mean-field approximation. In this paper, we have attempted to investigate whether beyond-mean-field effects on neutrino-neutrino coherent forward scattering could introduce flavor instabilities that are not otherwise foreseen. To this end, we have focused on the regime where neutrinos are degenerate and introduced a heuristic prescription for Pauli-blocking terms in the neutrino self-interaction Hamiltonian. Such corrections define an effective ELN distribution, which modifies the (anti)neutrino dispersion relation. Hence, ensembles that are on the verge of stability in the mean-field case can become unstable when the beyond-mean-field terms are included and vice versa.
Relying on the gyroscopic pendulum analogy introduced for fast flavor conversion [16, 21], we have investigated how the non-linear regime of the flavor evolution is modified by our beyond-mean-field prescription. Within the mean-field approximation, each (anti)neutrino polarization vector precesses around the potential generated by all other neutrinos in the ensemble. In the beyond-mean-field scenario, additionally, the length of the polarization vectors changes due to dynamical decoherence. The terms that cause this effect introduce an energy dependence in the collective potential and break the symmetry between the neutrino and the antineutrino equations of motion. As a result, flavor instabilities cascade down to small angular scales, and the evolution of the system is not periodic anymore.
Qualitatively, the outcome of flavor conversion including the beyond-mean-field effects in our toy model is similar to what is observed when incoherent scattering and vacuum mixing are included in the mean-field equations of motion, although the physics linked to each of these phenomena and their interplay with the self-interaction term in the Hamiltonian are different. We have neglected these additional terms in the equations of motion throughout this work, to highlight the impact of beyond-mean-field corrections on fast instabilities. However, collisions, vacuum mixing, and potential beyond-mean-field corrections (due to neutrino degeneracy) should coexist in the core of an astrophysical system.
@@ Acknowledgements.
Appendix A Correlation functions and scattering amplitudes
In neutrino-dense astrophysical sources, neutrino-neutrino scattering can be modeled as a Fermi interaction, as shown in the left panel of Fig. 9. The external lines represent the neutrino field $\nu_{\alpha}(\mathbf{p})=a_{\mathbf{p}}u_{\mathbf{p}}e^{-i\mathbf{p}\cdot\mathbf{r}}$, where $u_{\mathbf{p}}$ is the left-handed particle spinor. The amplitude of the process $\nu_{\alpha}(\mathbf{p})+\nu_{\beta}(\mathbf{q})\to\nu_{\alpha}(\mathbf{p^{\prime}})+\nu_{\beta}(\mathbf{q^{\prime}})$ can be written as:
$\begin{array}[]{rl}H_{\alpha\beta}(\mathbf{p^{\prime}},\mathbf{q^{\prime}};\mathbf{p},\mathbf{q})&=\mu[\bar{u}_{\mathbf{p^{\prime}}}\gamma^{\mu}P_{L}u_{\mathbf{p}}][\bar{u}_{\mathbf{q^{\prime}}}\gamma_{\mu}P_{L}u_{\mathbf{q}}]\\ &\times a^{\dagger}_{\alpha,\mathbf{p^{\prime}}}a_{\alpha,\mathbf{p}}a^{\dagger}_{\beta,\mathbf{q^{\prime}}}a_{\beta,\mathbf{q}}\,,\end{array}$ (14)
where $P_{L}\equiv(1-\gamma_{5})/2$. In dense environments, the two-particle scattering processes occur in the presence of other neutrinos that act as “spectators” (cf. the right panel of Fig. 9). The latter do not participate in the interaction; since particle spinors are normalized such that $\bar{u}_{\mathbf{p}}u_{\mathbf{p}}=1$, their contribution to the diagram is:
$H^{(3)}_{\alpha\beta\gamma}(\mathbf{p^{\prime}},\mathbf{q^{\prime}},\mathbf{k};\mathbf{p},\mathbf{q},\mathbf{k})=H_{\alpha\beta}(\mathbf{p^{\prime}},\mathbf{q^{\prime}};\mathbf{p},\mathbf{q})a^{\dagger}_{\gamma,\mathbf{k}}a_{\gamma,\mathbf{k}}\,.$ (15)
In this paper, we aim to investigate whether, in degenerate environments, this additional pair of creation/annihilation operators can have important effects on the development of flavor instabilies. At the mean-field level, the evolution of the single-particle density matrix in Eq. (1) is given by:
$i\dot{\rho}_{\mathbf{p},\gamma\kappa}=\langle[a^{\dagger}_{\kappa,\mathbf{p}}a_{\gamma,\mathbf{p}},H]\rangle\,,$ (16)
where brackets denote an ensemble average. To describe the effect of neutrino-neutrino scattering, the Hamiltonian $H$ must be of the form shown in Eq. (14), integrated over $\mathbf{p^{\prime}}$, $\mathbf{q^{\prime}}$ and $\mathbf{q}$. Wick’s theorem can be used to decompose the right-hand side of the equation into products of two-operator correlation functions. For example, for $\kappa=\gamma=\alpha=e$ and $\beta=x$, this yields terms such as:
$\begin{array}[]{rl}\langle a^{\dagger}_{e,\mathbf{p^{\prime}}}a_{e,\mathbf{p}}a^{\dagger}_{x,\mathbf{q^{\prime}}}a_{x,\mathbf{q}}a^{\dagger}_{e,\mathbf{p}}a_{e,\mathbf{p}}\rangle&=\langle a^{\dagger}_{e,\mathbf{p^{\prime}}}a_{e,\mathbf{p}}a^{\dagger}_{x,\mathbf{q^{\prime}}}a_{x,\mathbf{q}}\rangle\\ &=\langle a^{\dagger}_{e,\mathbf{p^{\prime}}}a_{e,\mathbf{p}}\rangle\langle a^{\dagger}_{x,\mathbf{q^{\prime}}}a_{x,\mathbf{q}}\rangle\\ &+\langle a^{\dagger}_{e,\mathbf{p^{\prime}}}a_{x,\mathbf{p}}\rangle\langle a_{e,\mathbf{p}}a^{\dagger}_{x,\mathbf{q^{\prime}}}\rangle\,,\end{array}$ (17)
where the first equality is given by the anti-commutation relation $a^{\dagger}_{e,\mathbf{p}}a_{e,\mathbf{p}}=1-a_{e,\mathbf{p}}a^{\dagger}_{e,\mathbf{p}}$. Since the mean-field approximation forbids correlations between different momentum modes, i.e. $\langle a^{\dagger}_{\alpha,\mathbf{p^{\prime}}}a_{\beta,\mathbf{p}}\rangle=(2\pi)^{3}\delta(\mathbf{p^{\prime}}-\mathbf{p})\rho_{\mathbf{p},\beta\alpha}$, all contributions vanish unless $\mathbf{p^{\prime}}=\mathbf{q}$ and $\mathbf{q^{\prime}}=\mathbf{p}$ (“forward scattering”) or $\mathbf{p^{\prime}}=\mathbf{p}$ and $\mathbf{q^{\prime}}=\mathbf{q}$ (no scattering). Thus, Eq. (16) provides a closed system of differential equations, describing neutrinos with fixed momenta propagating in an external potential. However, this approximation relies on the assumption that the neutrino gas is homogeneous, which is not an accurate picture of neutrino ensembles in dense astrophysical systems. Momentum correlations of the form $\langle a^{\dagger}_{\beta,\mathbf{p}}a_{\alpha,\mathbf{q}}\rangle$ must exist in these environments, e.g. due to spatial inhomogeneities. It is not clear whether two-particle scattering amplitudes are enough to model quantum kinetics. Indeed, applying Wick’s theorem to the ensemble average $\langle H^{(3)}_{exe}\rangle$ yields:
$\begin{array}[]{rl}\langle H^{(3)}_{exe}\rangle&=\langle H_{ex}\rangle\langle a^{\dagger}_{e,\mathbf{k}}a_{e,\mathbf{k}}\rangle-\langle a^{\dagger}_{e,\mathbf{q^{\prime}}}a_{e,\mathbf{k}}\rangle(\langle a^{\dagger}_{x,\mathbf{p^{\prime}}}a_{x,\mathbf{q}}\rangle\\ &\times\langle a^{\dagger}_{e,\mathbf{k}}a_{e,\mathbf{q}}\rangle+\langle a^{\dagger}_{x,\mathbf{p^{\prime}}}a_{e,\mathbf{p}}\rangle\langle a^{\dagger}_{e,\mathbf{k}}a_{x,\mathbf{p}}\rangle)+\langle a^{\dagger}_{x,\mathbf{p^{\prime}}}a_{e,\mathbf{k}}\rangle\\ &\times(\langle a_{x,\mathbf{p}}a^{\dagger}_{e,\mathbf{q^{\prime}}}\rangle\langle a^{\dagger}_{e,\mathbf{k}}a_{e,\mathbf{q}}\rangle+\langle a^{\dagger}_{e,\mathbf{q^{\prime}}}a_{e,\mathbf{p}}\rangle\langle a^{\dagger}_{e,\mathbf{k}}a_{x,\mathbf{p}}\rangle)\,.\end{array}$ (18)
The additional terms provided by the neutrino acting as a spectator may affect the two-particle result. Let us consider $\alpha=e$ and $\beta=\gamma=x$, i.e., $\nu_{e}(\mathbf{p})+\nu_{x}(\mathbf{q})+\nu_{x}(\mathbf{k})\to\nu_{e}(\mathbf{p^{\prime}})+\nu_{x}(\mathbf{q^{\prime}})+\nu_{x}(\mathbf{k})$. Processes of this form are Pauli-blocked if $\mathbf{q^{\prime}}=\mathbf{k}$. As argued in Sec. II.2, these interactions can be coherently enhanced if momentum transfer is small. Hence, neutrino refraction should be sensitive to the presence of degenerate spectator particles. The scenario sketched in Fig. 2 corresponds to $\alpha=e$, $\beta=\gamma=x$ and $\mathbf{p^{\prime}}=\mathbf{k}$. This process is allowed by the exclusion principle, but it cannot increase the flavor coherence of the incoming $\nu_{e}(\mathbf{p})$. In a physical setting, the outgoing particle would be in flavor superposition, and thus indistinguishable from the spectator, so the final state of the interaction must be anti-symmetric in these two particles. If their final momenta are also identical ($\mathbf{p^{\prime}}\sim\mathbf{k}$), the only possibility is that they become a singlet, so that they always occupy a different flavor eigenstate.
To account for this effect in a kinetic approach, it would be necessary to calculate the correlations $\langle a^{\dagger}_{e,\mathbf{p}}a_{x,\mathbf{q}}\rangle$ for all $\mathbf{p}$ and $\mathbf{q}$. This would require a self-consistent way to solve the higher-order equations in the BBGKY hierarchy, which is beyond the scope of this work. Instead, we assess whether new features could be introduced in flavor conversion by momentum correlations with spectator particles. For this purpose, we assume that they are responsible for a correction proportional to the neutrino occupation number in the refractive term of the quantum kinetic equation.
Figure 9: Left: Two-particle neutrino-neutrino interaction in the low-energy approximation. Right: Same as left panel, but in the presence of a spectator neutrino $\nu_{\gamma}(\mathbf{k})$. These diagrams can describe the same processes sketched in Fig. 2, setting $a=e$, $\gamma=b=x$.